OpenStudy (anonymous):

7^x-4 = 7 Is this problem going to be zero?

OpenStudy (lucaz):

7^(x-4) or 7^x-4 ?

OpenStudy (anonymous):

Solve and Check. 7^x-4 = 7

OpenStudy (zzr0ck3r):

\(7^{x-4}=7\implies x=5\)

OpenStudy (anonymous):

Oh, ok. 7 ^0 that's 1 correct?

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

\(7^1=7\)

OpenStudy (anonymous):

Yes, thank you...it's been a long time...thank you!